Charlene is the owner of Charlene’s Computer Care (3C) and has been in the business of servicing computers and computer systems at affordable prices for just over five years. 3C‘s primary clients are local businesses but include some residential customers. Their advertised approach is to analyze the problem, develop a range of IT options, explain their specific recommendation to the customer, and finally estimate the time and costs associated with the service prior to commencing work.

Their approach has proven successful, but competition is getting much stiffer, and Charlene is concerned that her approach to determining costs and prices for any particular job is not correctly incorporating all of the company's costs. She has heard about the concept of cost allocation and asks her accountant to create an allocation system for the business.

Her accountant decides to implement a single-driver allocation system and needs to develop a reliable cost function to estimate total indirect costs. She will use this estimate for the numerator of the single driver "overhead" rate, and she will use service hours as the cost driver and denominator in the "overhead" rate.

To estimate the parameters of this cost function, the accountant first studied the behavior of last month's costs, when the company spent 316 service hours repairing 39 computer systems. Here are the results of her analysis:

Cost Item

Fixed Cost

Variable Cost

  Utilities

$208

$75

  Office rent

$3,610

$0

  Administration

$3,134

$192

  Supplies

$0

$3,492

  Training

$370

$346

Total

$7,322

$4,105

Concerned that last month may not have been a typical month and her breakdown of fixed and variable costs not completely correct, the accountant decided to use the high-low method to create another cost function. She gathered individual cost and service hour data from the last six months; the following were for the two months when costs and service hours were the lowest and the highest:

Cost Item

low month

high month

  Utilities

$212

$305

  Office rent

$2,994

$2,994

  Administration

$3,132

$3,524

  Supplies

$1,519

$3,513

  Training

$605

$745

Total

$8,462

$11,081

  Service hours

247

573

  Customers

31

73

Next month, Charlene expects the company to spend 430 hours solving clients' computer system problems.


REQUIRED [ROUND ALL PER-UNIT COSTS TO TWO DECIMAL PLACES AND TOTAL COSTS TO THE NEAREST DOLLAR.]

Part A (6 tries; 5 points)
What is the accountant's estimate of total "overhead" next month if she uses the cost function based on her analysis of last month's cost data?   

Tries 0/6


Part B (6 tries; 5 points)
What is the accountant's estimate of total "overhead" next month if she uses the cost function based on the high-low method?   
Answer #1

Variable overhead per hour = 4105/316 = $12.99 per hour

Total overhead based on last month = Fixed overhead + variable overhead

= 7322 + 12.99*430

= $12,907.7

I.e. $12,908

b.Variable cost per unit as per high low method = Difference in cost/Difference in hours

= (11081-8462)/(573-247)

= $8.03 per hour

Fixed cost = 8462 – 247*8.03 = $6478.59

Overhead next month = 6478.59 + 8.03*430

= $9931.49